### v1

# PSGait: Multimodal Gait Recognition using Parsing Skeleton

Abstract
Gait recognition has emerged as a robust biometric modality due to its non-intrusive nature and resilience to occlusion. Conventional gait recognition methods typically rely on silhouettes or skeletons. Despite their success in gait recognition for controlled laboratory environments, they usually fail in real-world scenarios due to their limited information entropy for gait representations. To achieve accurate gait recognition in the wild, we propose a novel gait representation, named Parsing Skeleton. This representation innovatively introduces the skeleton-guided human parsing method to capture fine-grained body dynamics, so they have much higher information entropy to encode the shapes and dynamics of fine-grained human parts during walking. Moreover, to effectively explore the capability of the parsing skeleton representation, we propose a novel parsing skeleton-based gait recognition framework, named PSGait, which takes parsing skeletons and silhouettes as input. By fusing these two modalities, the resulting image sequences are fed into gait recognition models for enhanced individual differentiation. We conduct comprehensive benchmarks on various datasets to evaluate our model. PSGait outperforms existing state-of-the-art multimodal methods. Furthermore, as a plug-and-play method, PSGait leads to a maximum improvement of 10.9% in Rank-1 accuracy across various gait recognition models. These results demonstrate the effectiveness and versatility of parsing skeletons for gait recognition in the wild, establishing PSGait as a new state-of-the-art approach for multimodal gait recognition.
Index Terms:
Gait Recognition, Multimodal Fusion, Gait Representation, Human Parsing
I Introduction
Gait recognition aims to identify the target person by the motion pattern of human walking. Compared with face and fingerprint, gait has the advantages of being remotely accessible, non-contacting, and hard to disguise, which makes it a promising biometric for applications like ID authentication and social security[1, 2].
Existing gait representations for gait recognition primarily rely on silhouette sequences [3], Gait Energy Images (GEIs)[4], 2D skeletons[5], 3D skeletons[6], etc. Although these representations perform well in laboratory environments, they often fail in unconstrained situations, i.e., gait recognition in the wild [7, 8]. This failure is attributed to the low information entropy of binary silhouettes or 2D/3D skeletons, which are insufficient for effectively encoding the shapes and dynamics of the human body during walking. Recent research has increasingly focused on multimodal fusion methods in gait recognition[9, 10, 8, 11, 12], combining silhouettes and skeletons to create a more comprehensive representation of human gait. For example, Skeletongait++[9] utilized the Skeleton Map to represent the skeleton and integrated it with the silhouette using an attention mechanism. Although this approach has shown improvements, it relied on Gaussian points to represent the skeleton, neglecting detailed information about the body parts between these points. In gait recognition, the motion dynamics of different body parts vary significantly, and this disparity provides valuable complementary information for distinguishing individuals. Therefore, it is crucial to explore new representations with higher information entropy for gait recognition in the wild. Recent studies have demonstrated that human parsing provides a fine-grained and more comprehensive representation of gait, significantly enhancing the performance of gait recognition models[13, 14, 15, 16]. Although human parsing has shown promising results in gait recognition, conventional methods typically rely on RGB images, which are highly sensitive to factors like clothing variations and environmental occlusion, thereby reducing the accuracy of gait recognition in the wild.
Figure 1: The comparison of various gait representations. Only a single frame is displayed for brevity. Result shows the average Rank-1 on CCPG[17] for different representation inputs to GaitBase[18].
To address these issues, we propose a novel gait representation named Parsing Skeleton, which can be easily extracted through image processing methods based on the skeleton. Compared to existing gait recognition representations, the parsing skeleton offers three key advantages: 1) it contains higher information entropy at the pixel level, encoding more details about human motion; 2) it represents part-level knowledge, enabling models to learn fine-grained dynamics of different body parts during walking; 3) being derived from skeleton data, it improves the robustness and accuracy of gait recognition in real-world scenarios, outperforming existing RGB-based parsing methods. To effectively explore the capability of the parsing skeleton, we propose a parsing skeleton-based gait recognition framework named PSGait, which takes parsing skeletons and silhouettes as input. By fusing these two modalities through overlaying, the resulting image sequences are fed into gait recognition models for individual differentiation. We conduct comprehensive benchmarks on various datasets to evaluate our method against existing gait recognition approaches. PSGait achieves a maximum improvement of 5.2% in Rank-1 accuracy compared to the state-of-the-art multimodal method. Furthermore, replacing silhouettes with the fusion of silhouettes and parsing skeletons maximally leads to a 10.9% improvement in Rank-1 accuracy across various gait recognition models, underscoring the significant contribution of the plug-and-play parsing skeleton to gait recognition in real-world scenarios. These results validate the effectiveness and versatility of the parsing skeleton, establishing PSGait as a new state-of-the-art approach for multimodal gait recognition.
Figure 2: Overview of the PSGait framework. During data preprocessing, the silhouettes and skeletons are generated from video frames. In multimodal fusion, after the parsing section, the parsing skeletons are fused with silhouettes. The fused data is then fed into the gait recognition model for individual differentiation. The parsing section refines body part representations through selected points, creating circular and linear representations to enhance the robustness and accuracy of gait recognition in the wild. Both the data preprocessing and gait recognition models are interchangeable.
In summary, the contributions of this paper are as follows:
  • To the best of our knowledge, this is the first work to explore a simple yet effective gait representation for gait recognition, namely Parsing Skeleton, which demonstrates higher information entropy compared to existing silhouette- or skeleton-based representations.
  • We introduce PSGait, a novel gait recognition framework based on the parsing skeleton, which extracts comprehensive gait features with parsing skeletons and silhouettes, thereby enhancing the feature robustness in the wild.
  • Extensive experiments across various datasets validate the effectiveness and versatility of the parsing skeleton, establishing PSGait as a new state-of-the-art method for multimodal gait recognition.
II Related Work
Multimodal gait recognition aims to fuse multiple modalities to create a more comprehensive representation of gait. Specifically, SMPLGait[8] leveraged 3D geometric information from the SMPL model to enhance the gait appearance feature learning. BiFusion[10] integrated silhouettes and skeletons, using a graph convolutional network (GCN) with a multi-scale gait graph (MSGG) to extract complementary features from both modalities. Skeletongait++[9] employed the Skeleton Map to represent the skeleton and integrated it with the silhouette using an attention mechanism.
Human parsing in gait recognition typically relies on representations that contain fine-grained semantic information about individual body parts. These methods achieve simple but efficient human body parsing based on RGB images. Specifically, ParsingGait[13] combined CNNs and GCNs to effectively capture both global and local gait features, thereby enhancing gait recognition performance in real-world scenarios. GaitParsing[14] decomposed gait into specific body parts using human semantic parsing and employed a dual-branch feature extraction network to process both holistic and part-level features. LandmarkGait[16] introduced an unsupervised method that converts silhouettes into consistent landmarks for fine-grained parsing, utilizing a multi-scale feature extractor to capture both global and part-level features for gait recognition.
III Method
The PSGait proposed in this study is illustrated in Fig. 2. First, during data preprocessing, the silhouette and skeleton for each frame of the gait video sequence are obtained through segmentation and pose estimation. Next, after the parsing section, the Parsing Skeletons are fused with silhouettes to obtain the multimodal gait representation. Finally, this multimodal gait representation, which encapsulates comprehensive information, is fed into the gait recognition model to enhance its classification performance.
III-A Parsing Section
The main objective of this section is to obtain representations with fine-grained semantic information of various body parts from the skeleton. First, the point selection module involves two steps: data cleaning and human body part selection. To mitigate the impact of erroneous results from the pose estimation during data preprocessing, a confidence threshold $\tau$ is set to filter out unreliable points. Only points with confidence $c_{i}$ exceeding the threshold $\tau$ are used. Additionally, any points outside the valid region ($H\times W$) are disregarded. Therefore, a valid point ($x_{i},y_{i}$) must satisfy the following condition:
[…]
According to the joint point configuration in COCO17, we selected the corresponding joint points for each body part, including the head and limbs. For the head, three points were selected: the nose and both ears. For each limb, the joint points at both ends were chosen. This results in the following mapping between human body joints $k_{i}$ and their corresponding body parts $P_{j}$:
$h(P_{j})=\{k_{1},k_{2},\dots,k_{m}\}$ (2)
When rendering, we need to ensure that the parsing skeleton’s information is effectively captured by the CNN when it is fed into the gait recognition model with uniform size. Additionally, different colors $c$ are assigned to each body part to enhance the fine-grained information of the human body. For the head representation, we introduce circles centered at the three key head points ($x,y$), with a specified radius $r$, encompassing the around the pixel $(u,v)$:
$\text{circle}(x,y,r,c)=\{(u,v)\mid(u-x)^{2}+(v-y)^{2}\leq r^{2}\}$ (3)
For the representation of each limb, a line of specified width is parsed on the image:
$\text{line}((x_{1},y_{1}),(x_{2},y_{2}),w,c)=\left\{(u,v)\mid d\leq\frac{w}{2}\right\}$ (4)
where $d$ represents the distance of the pixel $(u,v)$ from the line, $c$ denotes the color assigned to this part, and $w$ is the width of this line.
III-B Fusion Strategy
In this study, we investigate two distinct fusion strategies for combining parsing skeleton and silhouette in representation learning. The objective of the fusion is to generate an effective input that captures both the global human shape from the silhouette and the fine-grained local structural relationships from the parsing skeleton.
Overlay-Based Fusion: The first strategy is a straightforward overlay operation, where the parsing skeleton representation (lines and circles) is directly superimposed onto the silhouette image $M(x,y)$. Specifically, the fused image $F_{\text{overlay}}(x,y)$ is defined as follows:
$F_{\text{overlay}}(x,y)=\max\left(M(x,y),S(x,y)\right),$ (5)
where $M(x,y)$ is the binary silhouette image, and $S(x,y)$ represents the parsing skeleton. The $\max$ operator ensures that the parsing skeleton structure is not overwritten by the silhouette but is integrated naturally. The overlay method produces an RGB fused image, where the silhouette and parsing skeleton share the same spatial space, allowing the network to jointly learn the global shape features and structural details with minimal processing complexity.
Multi-Channel Fusion: In the second strategy, we introduce a more granular representation by separating each parsing skeleton component (lines and circles) into independent channels. Specifically, each limb and head point are encoded in separate channels, and the silhouette $M(x,y)$ is placed in the final channel. This results in a multi-channel representation $F_{\text{multi}}$, where the fused feature map has 12 channels:
$F_{\text{multi}}(x,y,c)=\begin{cases}S_{c}(x,y),&\text{if }c=1,2,\dots,11\\ M(x,y),&\text{if }c=12\end{cases}$ (6)
where $S_{c}(x,y)$ represents the $c$-th parsing skeleton component (either a line or a circle), and the 12th channel contains the silhouette image. While this strategy provides a more explicit and structured representation, enabling the network to focus on individual components of the parsing skeleton and silhouette, it inherently increases the complexity of the input and may dilute spatial relationships across channels.
III-C Align Center
After the data fusion stage, the center alignment operation is performed to ensure that the human body contour and the resolved skeleton are centered in the image under the premise of uniform input image size. This alignment facilitates subsequent model processing and feature extraction. Given a binary silhouette image $M(x,y)$, where $M(x,y)\in\{0,1\}$, the upper bound $y_{\text{top}}$ and lower bound $y_{\text{btm}}$ of the human body region are calculated as follows:
$y_{\text{top}}=\min\left\{y\,\middle|\,\sum_{x}M(x,y)>0\right\}.$ (7)
$y_{\text{btm}}=\max\left\{y\,\middle|\,\sum_{x}M(x,y)>0\right\}.$ (8)
Next, the image is cropped based on the upper and lower bounds of the human region and resized to the target height while preserving the aspect ratio, resulting in the image $M^{\prime}(x,y)$. The center of the horizontal axis is calculated as follows:
$x_{c}=\frac{\sum_{x,y}x\cdot M^{\prime}(x,y)}{\sum_{x,y}M^{\prime}(x,y)}$ (9)
Using the horizontal axis center $x_{c}$, the image is cropped to a target size of $H^{\prime}\times W^{\prime}$.
TABLE I: Quantitative comparison of gait recognition methods on the SUSTech1K[19] is presented. The best performances are indicated in bold, and the second-best methods are underlined. †We modified Skeletongait++ to match the settings used in GaitBase.
Input Method Probe Sequence(R-1) Overall
Normal Bag Clothing Carrying Umbrella Uniform Occlusion Night R-1 R-5
Skeleton GaitGraph2(CVPRW2022)[20] 22.2 18.2 6.8 18.6 13.4 19.2 27.3 16.4 18.6 40.2
GaitTR(ES2023)[21] 33.3 31.5 21.0 30.4 22.7 34.6 44.9 23.5 30.8 56.0
Silhouette GaitSet(AAAI2019)[22] 69.1 68.2 37.4 65.0 63.1 61.0 67.2 23.0 65.0 84.8
GaitPart(CVPR2019)[23] 62.2 62.8 33.1 59.5 57.2 54.8 57.2 21.7 59.2 80.8
GaitGL(ICCV2021)[24] 67.1 66.2 35.9 63.3 61.6 58.1 66.6 17.9 63.1 82.8
GaitBase(CVPR2023)[18] 81.5 77.5 49.6 75.8 75.5 76.7 81.4 25.9 76.1 89.4
DeepGaitV2(Arxiv2023)[25] 83.5 79.5 46.3 76.8 79.1 78.5 81.1 27.3 77.4 90.2
[…]
Skeleton+Silhouette BiFusion(MTA2023)[10] 69.8 62.3 45.4 60.9 54.3 63.5 77.8 33.7 62.1 83.4
SkeletonGait++†(AAAI2024)[9] 84.7 82.4 46.7 81.8 81.2 82.6 85.1 44.3 81.2 93.2
[…]
TABLE II: Quantitative comparison of gait recognition methods is conducted across two authoritative datasets, CCPG[17] and Gait3D[8]. †We modified Skeletongait++ to match the settings used in GaitBase.
Input Method Testing Datasets
CCPG Gait3D
CL UP DN BG Mean Rank-1 Rank-5 mAP mINP
Skeleton GaitGraph2(CVPRW2022)[20] 5.0 5.3 5.8 6.2 5.6 11.1 - - -
GaitTR(ES2023)[21] 15.7 18.3 18.5 17.5 17.5 6.6 - - -
Silhouette GaitSet(AAAI2019)[22] 60.2 65.2 65.1 68.5 64.8 36.7 58.3 30.0 17.3
GaitPart(CVPR2020)[23] 64.3 67.8 68.6 71.7 68.1 28.2 47.6 21.6 12.4
GaitBase(CVPR2023)[18] 71.6 75.0 76.8 78.6 75.5 64.6 - - -
[…]
Skeleton+Silhouette SkeletonGait++†(AAAI2024)[9] 77.0 81.4 83.0 88.8 82.6 64.8 80.5 54.8 36.1
[…]
IV Experience
IV-A Datasets
We evaluated our proposed method on three commonly used datasets: SUSTech1K[19] and CCPG[17], which are collected in fully constrained laboratory environments, and Gait3D[8], which is collected from real-world scenarios.
SUSTech1K[19] contains gait data from 1,050 subjects, covering 8 different conditions: normal (NM), bag (BG), clothes changing (CL), carrying (CR), umbrella (UB), uniform (UF), occlusion (OC), and night (NT), as well as 12 viewpoints (ranging from 0° to 315°). The diversity and granularity of the dataset make SUSTech1K[19] an ideal platform for studying the performance of gait recognition in complex environments.
CCPG[17] was designed to address the challenge of gait recognition under clothing change conditions and is particularly suitable for dynamic person recognition in video surveillance. The dataset includes 200 identities and more than 16,000 sequences, with each identity captured from 10 viewpoints under seven different clothing change conditions, covering both indoor and outdoor environments. It provides two data modalities: RGB images and silhouette images.
Gait3D[8] is a large-scale gait dataset collected in real-world environments. It was recorded in a supermarket and contains 1,090 hours of video footage with a resolution of 1,920×1,080 and a frame rate of 25 FPS. Gait3D provides four modalities: silhouettes, 2D and 3D skeletons, and 3D meshes.
TABLE III: Impact of circle radius and line width on the SUSTech1K[19] dataset using GaitBase[18].
Circle Radius Line Width Probe Sequence(R-1) Overall
Normal Bag Clothing Carrying Umbrella Uniform Occlusion Night R-1 R-5
3 3 82.1 80.2 47.0 77.9 80.8 79.4 87.1 33.8 78.8 91.7
10 12 86.1 84.5 55.0 82.9 85.4 84.1 90.7 47.1 83.4 94.3
20 24 84.6 83.1 51.3 80.8 83.0 81.2 90.0 43.1 81.6 93.7
TABLE IV: Impact of input modalities and fusion strategies on the SUSTech1K[19] dataset using GaitBase[18].
Skeleton Silehouette Overlay Probe Sequence(R-1) Overall
Normal Bag Clothing Carrying Umbrella Uniform Occlusion Night R-1 R-5
52.5 50.1 35.2 48.4 40.3 51.3 61.1 22.0 50.3 72.5
81.5 77.5 49.6 75.8 75.5 76.7 81.4 25.9 76.1 89.4
82.7 81.2 47.8 78.9 81.3 79.8 88.2 35.1 79.7 92.2
86.1 84.5 55.0 82.9 85.4 84.1 90.7 47.1 83.4 94.3
IV-B Implementation Details
For data preprocessing, the original data is used to extract silhouette information from all three datasets. Since skeleton data is not available for the CCPG dataset, we utilize the Sapiens[26] method for pose estimation. The original skeleton data is used for the other two datasets, and the format of points follows COCO17. A confidence threshold of 0.1 is applied.
In the parsing stage, to ensure that the Parsing Skeleton features of the processed image can be effectively captured by the CNN, the radius of the circle representing the head and the width of the line segment representing the limbs are determined based on the image size ratio in each dataset. Specifically, for the SUSTech1K[19] dataset, a radius of 10 and a line width of 12 are used. For the CCPG[17] dataset, the radius and line width are set to 16 and 20, respectively. Similarly, for the Gait3D[8] dataset, the radius and line width are also set to 16 and 20, respectively. The colors selected for parsing each body part are shown in Fig. 2.
To evaluate the versatility of our method, we apply it to two state-of-the-art gait recognition models: GaitBase[18] and DeepGaitV2[25]. The size of the fused image is adjusted to $64\times 44$ before being input into the gait recognition model. For model training, we largely follow the parameter settings from OpenGait[18], an excellent open-source framework in the field of gait recognition. Specifically, for the SUSTech1K[19] dataset, the batch size is set to $(8,8)$, and learning rate milestones are scheduled at 20k, 30k, and 40k steps, with a total of 50k training steps. For the CCPG[17] dataset, we use a batch size of $(8,16)$ with learning rate milestones at 20k, 40k, and 50k steps, and a total of 60k iterations. For the Gait3D[8] dataset, the batch size is set to $(32,4)$, with learning rate milestones occurring at 20k, 40k, and 50k steps, and a total of 60k iterations. The optimization process is configured using the SGD solver. Specifically, the learning rate $lr$ is initialized at 0.1, providing an appropriate step size for updating the model weights during training. A momentum value of 0.9 is employed to accelerate convergence by reducing oscillations and incorporating a weighted history of previous gradients.
IV-C Comparison with State-of-the-art Methods
As shown in Table I and Table II, we evaluate the performance on three datasets. The results highlight the strong competitiveness and versatility of our method.
Competitiveness. We compare and validate the advantages of our method across three datasets. On the SUSTech1K[19] dataset (Table I), compared to SkeletonGait++, Ours+GaitBase improves the Rank-1 accuracy from 81.2% to 83.4% (+2.2%). This demonstrates the effectiveness of our method under the same baseline settings. Furthermore, Ours+DeepGaitV2 achieves the best overall performance on SUSTech1K, attaining a Rank-1 accuracy of 87.9% and a Rank-5 accuracy of 94.8%. On the CCPG[17] dataset (Table II), compared to SkeletonGait++[9], Ours+GaitBase improves the mean accuracy from 82.6% to 86.4% (+3.8%). On the Gait3D[8] dataset (Table II), Rank-1 accuracy is improved from 64.8% to 70.0% (+5.2%). Additionally, Ours+GaitBase and Ours+DeepGaitV2 achieve the best overall performance on CCPG and Gait3D, with a mean accuracy of 86.4% on CCPG and a Rank-1 accuracy of 85.6% and mAP of 69.4% on Gait3D, outperforming all other state-of-the-art methods. These improvements underscore the robustness and adaptability of our approach across diverse benchmarks and evaluation metrics.
Versatility. We further demonstrate the versatility of our approach by integrating it into two baseline models, GaitBase[18] and DeepGaitV2[25], and comparing the resulting performance improvements. On the three datasets mentioned above, our method provides a significant boost over the existing baselines. Specifically, on the SUSTech1K dataset[19], Ours+GaitBase improves the Rank-1 accuracy from 76.1% to 83.4% (+7.3%), while Ours+DeepGaitV2 enhances the Rank-1 accuracy from 77.4% to 84.6% (+7.2%). On the CCPG[17] dataset, integrating our method with GaitBase increases the mean accuracy from 75.5% to 86.4% (+10.9%). For DeepGaitV2, the mean accuracy improves from 83.3% to 85.6% (+2.3%). On Gait3D[8], Ours+GaitBase improves the Rank-1 accuracy from 64.6% to 70.0% (+5.4%). For DeepGaitV2, the Rank-1 accuracy increases from 74.4% to 76.2% (+1.8%), and the mAP rises from 65.8% to 69.4% (+3.6%). These consistent performance improvements across multiple datasets and baseline models highlight the strong versatility of our method. It effectively enhances the performance of any gait recognition model, making it plug-and-play and universally applicable to various scenarios and datasets.
IV-D Ablation Study
To validate the effectiveness of our method, we conducted ablation studies on the SUSTech1K[19] dataset using GaitBase[18] as the baseline gait recognition model. We conducted experiments to assess the impact of key parameters, including circle radius, line width, input modalities, and different data fusion strategies.
Influence of Circle Radius and Line Width. As shown in Table III, we evaluated the impact of circle radius and line width on recognition performance. A smaller setting results in a relatively low Rank-1 accuracy of 82.1%, while a moderate setting achieves the best Rank-1 accuracy of 86.1% and Rank-5 accuracy of 94.3%. However, further increasing the radius and width, slightly reduces Rank-1 accuracy to 84.6%, indicating that finding an optimal balance is crucial for performance. The results indicate that choosing an appropriate circle radius and line width is critical for effectively extracting parsing skeleton features. If these values are too small, the parsing skeleton cannot be effectively captured by the CNN after resizing the image. Conversely, if they are too large, the parsing skeleton may obscure the silhouette, impeding the extraction of human shape features by gait recognition models.
Impact of Data Fusion Strategies. We further analyzed the contribution of different modalities and fusion methods (parsing skeleton, silhouette, and overlay) to overall performance. As shown in Table IV, using parsing skeletons or silhouettes alone does not yield particularly high performance, indicating that relying on either type of data alone is insufficient for accurate gait recognition. Incorporating silhouette information significantly improves performance, with Rank-1 increasing to 79.7% and Rank-5 reaching 92.2%. In this case, the data fusion strategy employed is a multi-channel fusion method. When parsing skeleton and silhouette are overlaid, the model achieves the best performance, with a Rank-1 accuracy of 83.4% and a Rank-5 accuracy of 94.3%. The overlay fusion strategy capitalizes on the complementary strengths of skeleton and silhouette data, allowing the model to capture both global shape and local structural details. These results demonstrate that our fusion strategy is both effective and achieves an optimal balance between accuracy and robustness.
V Conclusion
In this paper, we propose a novel gait representation, Parsing Skeleton, which leverages skeleton-guided human parsing to capture fine-grained body dynamics. This representation addresses the limitations of traditional gait recognition methods that rely on silhouettes or skeletons with low information entropy, by encoding more detailed and comprehensive features of human motion. To effectively exploit the parsing skeleton, we propose PSGait, a new gait recognition framework that integrates parsing skeletons with silhouettes to enhance feature robustness and improve individual differentiation, particularly in unconstrained, real-world scenarios. Extensive experiments across various datasets demonstrate that PSGait outperforms existing state-of-the-art multimodal gait recognition methods. As a plug-and-play method, PSGait leads to significant performance improvements across various models. This work validates the effectiveness and versatility of the parsing skeleton representation for real-world gait recognition tasks, establishing PSGait as a new state-of-the-art approach for multimodal gait recognition.